Venugopalramana has kindly calculated and ploted the data from y=3cos(2t+ 40 degrees),0 < t < 360 degrees. But I do not know how the data was calculated and how he plotted the data. Could Venugopalramana or another tutor please explain the steps to calculating the data. Thankkyou
Answer by venugopalramana(3286)

sketch y=3cos(2t+ 40 degrees),0 < t < 360 degrees
CALCULATE THE DATA AS SHOWN BELOW AND PLOT THEM ON A GRAPH.IT WILL
LOOK AS GIVEN BELOW.
T-DEG. T-RADIANS y=3cos(2t+ 40 degrees)
0 0.00 2.30........OK I WILL SHOW ONE EXAMPLE.LET US DO FOR T=30 DEGREES
WE CAN CALCULATE WITH DEGREES IF YOU HAVE CALCULATOR WITH DEGREE
MEASURE.....FIND 2T+40=2*30+40=60+40=100...NOW FIND COS(100) DEGREES IN
CALCULATOR..YOU WILL GET IT AS -0.1733....MULTIPLY WITH 3 TO GET
-0.52...CONTINUE THIS WAY FOR DIFFERENT VALUES OF T STARTING FROM 0
DEGREES AND GOING IN STEPS OF 30 OR 60 DEGREES AS YOU LIKE.
AFTER TABULATING AS SHOWN HERE TAKE A SUITABLE SCALE FOR X AND Y AXIS
AND PLOT THEM FOR T=0 TO 360 DEGREES AS DESIRED....VENUGOPAL
30 0.52 -0.52
60 1.05 -2.82
90 1.57 -2.30
120 2.09 0.52
150 2.62 2.82
180 3.14 2.30
210 3.67 -0.52
240 4.19 -2.82
270 4.71 -2.30
300 5.24 0.52
330 5.76 2.82
360 6.28 2.30